OpenStudy (anonymous):

Suppose y varies directly as x. Write a direct variation that relates x and y. Then solve. 1. If y=6 when x=10, find x when y=18 2. If y=4 and 1/4 when x=3/4, find y when x=4 and 1/2

OpenStudy (anonymous):

y = k*x, 1. 6 = k*10 6/10 = k